### v2

@@ 3.2 Nearly Linear Models
Proposition 2.
Let $P_{x}(\cdot)\subset{\mathcal{P}}({\mathcal{P}}(\mathds{X}))$ denote the training distribution of the mean-field terms for the state $x\in\mathds{X}$. Let Assumption 6 hold. For the estimate models, $\hat{c}(x,u,\mu)$ and $\hat{\mathcal{T}}(\cdot|x,u,\mu)$ defined based on the least square method (see (3.2)), we have that
@@
Proof.
We first note that since the learned $\theta^{*}_{(x,u)}$ minimizes the $L_{2}$ distance to the true model under the training measure $P_{x}$, (6) implies that
@@
where we used $\|{\bf\Phi}_{(x,u)}(\mu)\|_{2}\leq\sqrt{d}$ since we assume that $\|\Phi^{j}_{(x,u)}(\cdot)\|_{\infty}\leq 1$ via Assumption 2.
For the proof of the error bound of the estimate kernel $\hat{\mathcal{T}}(\cdot|x,u,\mu)$ we follow identical steps, recalling that by construction ${\bf\Phi}_{(x,u)}^{\intercal}(\mu){\bf Q}^{*,j}{(x,u)}$ minimizes the $L_{2}$ distance to $\mathcal{T}(j|x,u,\mu)$ and using Assumption 6, we write that
@@
where we used $\sum_{j}\epsilon_{j}\leq\epsilon$ with Assumption 6. ∎
@@ 3.3 A Special Case: Linear Approximation via Discretization
The following then is an immediate result of Assumption 1.
Proposition 3.
Let $\theta_{x,u}=[\theta^{1}_{x,u},\dots,\theta^{d}_{x,u}]^{\intercal}$ and ${\bf Q}_{x,u}(\cdot)=\left[Q^{1}_{x,u}(\cdot),\dots,Q^{d}_{x,u}(\cdot)\right]^{\intercal}$ be given by (3.3). If the training measure $P(\cdot)$ is such that $P(B_{i})>0$ for each quantization bin $B_{i}$, under Assumption 1, we then have that
@@ 4 Error Analysis for Control with Misspecified Models
We will analyze the infinite population and the finite population settings separately. We note that some of the following results (e.g. Lemma 3) have been studied in the literature to establish the connection between the $N$-agent control problems and the limit mean-field control problem without the model mismatch aspect. That is, existing results study what happens if one uses the infinite population solution for the finite population control problem with perfectly known dynamics (see e.g. [33, 5, 32]). However, we present the proof of every result for completeness and because of the connections in the analysis we follow throughout the paper. Furthermore, the existing results are often stated under slightly different assumptions and settings such as being stated only for closed loop policies, or only for policies that are open loop in the sense that they are measurable with respect to the noise process.
@@ 4.1 Error Bounds for Infinitely Many Agents
As we have observed in Example 1.2, even when agents agree on the model knowledge, without coordination on which policy to follow, the optimality may not be achieved. Therefore, we assume that after the learning period, the team of agents collectively agrees on the cost and transition models given by $\hat{c}(x,u,\mu)$ and $\hat{\mathcal{T}}(\cdot|x,u,\mu)$ and designs policies for this model. We will assume that
@@
In what follows, previously introduced constants $K_{c},K_{f}$ and $\delta_{T}$ will be used often. We refer the reader to Assumption 1 for $K_{c},K_{f}$, and equation (6) for $\delta_{T}$.
@@ 4.1.1 Error Bounds for Closed Loop Control
Theorem 3.
Consider the closed loop policy $\hat{g}$ in (32) designed for an estimate model $\hat{\mathcal{T}},\hat{c}$ which satisfies (4.1) for the infinite population dynamics. Under Assumption 1, if $\beta K<1$
@@
Proof.
We start with the following upper-bound
@@ 4.1.2 Error Bounds for Open Loop Control
Theorem 4.
Consider the open loop policy $\hat{g}$ in (32) which is designed for an estimate model that satisfies (4.1) for the infinite population dynamics. Under Assumption 1, if $\beta K<1$,
@@
Proof.
We start with the following upper-bound
@@
This is the bound for the first term in (36), combining this with the upper-bound on the second term in (36) by Lemma 1, we can complete the proof. ∎
Lemma 1.
Under Assumption 1, if $\beta K<1$
@@
Proof.
The proof can be found in the appendix B. ∎
@@ 4.2.1 Error Bounds for Open Loop Control
Theorem 5.
Under Assumption 1, if each agent follows the steps summarized above, we then have that
@@
where we have used Lemma 2 which is presented below. ∎
Lemma 2.
Let $x_{t}^{i}$ be the state of the agent $i$ at time $t$ when each agent follows the open-loop policy $\hat{\gamma}(\cdot|x^{i}_{t},\hat{\mu}_{t})$ in an $N$-agent control dynamics. We denote by ${\bf x}_{t}$ the vector of the states of $N$ agents at time $t$. Under Assumption 1, we then have that
@@
Proof.
The proof can be found in Appendix C. ∎
Lemma 3.
Under Assumption 1,
@@
Proof.
The proof can be found in the appendix D. ∎
@@ 4.2.2 Error Bounds for Closed Loop Control
Theorem 6.
Under Assumption 1, if each agent follows the steps summarized above, we then have that
@@
Proof.
The proof follows very similar steps to the results we have proved earlier. For some $\hat{\mu}_{0}=\mu_{0}=\mu_{\bf x_{0}}=\mu^{N}$
@@ 5 Numerical Study
  • Furthermore, to encourage a balanced distribution, we penalize deviations from the $40$%-$60$% distribution by introducing a cost $10\times(\mu(0)-0.4)^{2}$ where $\mu(0)$ is the fraction of agents in the surge zones.
For the dynamics, we assume that a non-surge area has a fixed probability $0.2$ of becoming a surge area in the next time step. Furthermore, we assume that a surge area has a probability $0.20.7*\mu(0)+$ 0.2 of becoming a non-surge area, indicating that as more drivers there are in a surge area ($\mu(0)$ is high), the likelihood of it remaining a surge zone decreases (due to increased supply). This then defines the transition probabilities as follows:
$Pr(X_{t+1}^{i}=1|X_{t}^{i}=0,U_{t}^{i}=0,\mu)=0.7*\mu(0)+0.2$
@@
Near optimality of learned models and infinite population approximations. Figure 1 shows the value functions loss for different values $N$, the number of agents in the system. We graph the loss functions under 3 settings:
@@
For the loss, we compare the value of the policy with the optimal value in an infinite population environment. Furthermore, we assume that the initial distribution is $\mu_{0}=1/2\delta_{0}+1/2\delta_{1}$.
@@ 6 Conclusion
Appendix A Proof of Proposition 1
@@ Appendix B Proof of Lemma 1
Finally, a slight modification of [6, Lemma 6] for finite $\mathds{X},\mathds{U}$ can be used to show that

### v3

@@ 3.2 Nearly Linear Models
Proposition 3.1
Let $P_{x}(\cdot)\subset{\mathcal{P}}({\mathcal{P}}(\mathds{X}))$ denote the training distribution of the mean-field terms for the state $x\in\mathds{X}$. Let Assumption 3.1 hold. For the estimate models, $\hat{c}(x,u,\mu)$ and $\hat{\mathcal{T}}(\cdot|x,u,\mu)$ defined based on the least square method (see (3.2)), we have that
@@
Proof  We first note that since the learned $\theta^{*}_{(x,u)}$ minimizes the $L_{2}$ distance to the true model under the training measure $P_{x}$, (3.1) implies that
@@
where we used $\|{\bf\Phi}_{(x,u)}(\mu)\|_{2}\leq\sqrt{d}$ since we assume that $\|\Phi^{j}_{(x,u)}(\cdot)\|_{\infty}\leq 1$ via Assumption 2.1.
For the proof of the error bound of the estimate kernel $\hat{\mathcal{T}}(\cdot|x,u,\mu)$ we follow identical steps, recalling that by construction ${\bf\Phi}_{(x,u)}^{\intercal}(\mu){\bf Q}^{*,j}{(x,u)}$ minimizes the $L_{2}$ distance to $\mathcal{T}(j|x,u,\mu)$ and using Assumption 3.1, we write that
@@
where we used $\sum_{j}\epsilon_{j}\leq\epsilon$ with Assumption 3.1.
@@ 3.3 A Special Case: Linear Approximation via Discretization
The following then is an immediate result of Assumption 1.1.
Proposition 3.2
Let $\theta_{x,u}=[\theta^{1}_{x,u},\dots,\theta^{d}_{x,u}]^{\intercal}$ and ${\bf Q}_{x,u}(\cdot)=\left[Q^{1}_{x,u}(\cdot),\dots,Q^{d}_{x,u}(\cdot)\right]^{\intercal}$ be given by (3.3). If the training measure $P(\cdot)$ is such that $P(B_{i})>0$ for each quantization bin $B_{i}$, under Assumption 1.1, we then have that
@@ 4 Error Analysis for Control with Misspecified Models
We will analyze the infinite population and the finite population settings separately. We note that some of the following results (e.g. Lemma 16) have been studied in the literature to establish the connection between the $N$-agent control problems and the limit mean-field control problem without the model mismatch aspect. That is, existing results study what happens if one uses the infinite population solution for the finite population control problem with perfectly known dynamics (see e.g. Motte and Pham (2023); Bäuerle (2023); Motte and Pham (2022)). However, we present the proof of every result for completeness and because of the connections in the analysis we follow throughout the paper. Furthermore, the existing results are often stated under slightly different assumptions and settings such as being stated only for closed loop policies, or only for policies that are open loop in the sense that they are measurable with respect to the noise process.
@@ 4.1 Error Bounds for Infinitely Many Agents
As we have observed in Example 2, even when agents agree on the model knowledge, without coordination on which policy to follow, the optimality may not be achieved. Therefore, we assume that after the learning period, the team of agents collectively agrees on the cost and transition models given by $\hat{c}(x,u,\mu)$ and $\hat{\mathcal{T}}(\cdot|x,u,\mu)$ and designs policies for this model. We will assume that
@@
In what follows, previously introduced constants $K_{c},K_{f}$ and $\delta_{T}$ will be used often. We refer the reader to Assumption 1.1 for $K_{c},K_{f}$, and equation (6) for $\delta_{T}$.
@@ 4.1.1 Error Bounds for Closed Loop Control
Theorem 10
Consider the closed loop policy $\hat{g}$ in (32) designed for an estimate model $\hat{\mathcal{T}},\hat{c}$ which satisfies (4.1) for the infinite population dynamics. Under Assumption 1.1, if $\beta K<1$
@@
Proof  We start with the following upper-bound
@@ 4.1.2 Error Bounds for Open Loop Control
Theorem 11
Consider the open loop policy $\hat{g}$ in (32) which is designed for an estimate model that satisfies (4.1) for the infinite population dynamics. Under Assumption 1.1, if $\beta K<1$,
@@
Proof  We start with the following upper-bound
@@
This is the bound for the first term in (36), combining this with the upper-bound on the second term in (36) by Lemma 12, we can complete the proof.
Lemma 12
Under Assumption 1.1, if $\beta K<1$
@@
Proof  The proof can be found in the appendix B.
@@ 4.2.1 Error Bounds for Open Loop Control
Theorem 14
Under Assumption 1.1, if each agent follows the steps summarized above, we then have that
@@
where we have used Lemma 15 which is presented below.
Lemma 15
Let $x_{t}^{i}$ be the state of the agent $i$ at time $t$ when each agent follows the open-loop policy $\hat{\gamma}(\cdot|x^{i}_{t},\hat{\mu}_{t})$ in an $N$-agent control dynamics. We denote by ${\bf x}_{t}$ the vector of the states of $N$ agents at time $t$. Under Assumption 1.1, we then have that
@@
Proof  The proof can be found in Appendix C.
Lemma 16
Under Assumption 1.1,
@@
Proof  The proof can be found in the appendix D.
@@ 4.2.2 Error Bounds for Closed Loop Control
Theorem 17
Under Assumption 1.1, if each agent follows the steps summarized above, we then have that
@@
Proof  The proof follows very similar steps to the results we have proved earlier. For some $\hat{\mu}_{0}=\mu_{0}=\mu_{\bf x_{0}}=\mu^{N}$
@@ 5 Numerical Study
  • Furthermore, to encourage a balanced distribution, we penalize deviations from $40$%-$60$% distribution by introducing a cost $10\times(\mu(0)-0.4)^{2}$ where $\mu(0)$ is the fraction of agents in the surge zones.
For the dynamics, we assume that a non-surge area has a fixed probability $0.2$ of becoming a surge area in the next time step. Furthermore, we assume that a surge area has a probability $0.20.7\times\mu(0)+$ 0.2 of becoming a non-surge area, indicating that as more drivers there are in a surge area ($\mu(0)$ is high), the likelihood of it remaining a surge zone decreases (due to increased supply). This then defines the transition probabilities as follows:
$Pr(X_{t+1}^{i}=1|X_{t}^{i}=0,U_{t}^{i}=0,\mu)=0.7\times\mu(0)+0.2$
@@
Near optimality of learned models and infinite population approximations. Figure 1 shows the value loss for different values of the number of agents in the system. We graph the loss functions under 3 settings:
@@
For the loss, we compare the value of the learned approximate policy with the optimal value in an infinite population environment. Furthermore, we assume that the initial distribution is $\mu_{0}=1/2\delta_{0}+1/2\delta_{1}$.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
E. Bayraktar is partially supported by the National Science Foundation under grant DMS-2106556 and by the Susan M. Smith chair.
Appendix A Proof of Proposition 2.1
@@ Appendix B Proof of Lemma 12
Finally, a slight modification of (Bayraktar et al., 2025, Lemma 6) for finite $\mathds{X},\mathds{U}$ can be used to show that